### v1

@@ 6 Results
This is the same interface (and to some extent, the same code) as lean4checker [^7] https://github.com/leanprover/lean4checker, which does the exact same thing but calls addDecl instead, effectively using the C++ kernel as an external verifier for Lean proofs. It may seem odd to use the Lean kernel as an external verifier for itself, but this can catch cases where malicious (or confused) Lean code makes use of the very powerful metaprogramming framework to simply bypass the kernel and add constants to the environment without typechecking them.
@@
lean4export lean4lean ratio
@@
Std 32.49 s 45.74 s 1.40
Mathlib (+ Std + Lean) 44.54 min 58.79 min 1.32
Figure 2: Comparison of the original C++ implementation of the Lean kernel (lean4export) with lean4lean (Lean) on major Lean packages. Tests were performed on a 12 core 12th Gen Intel i7-1255U @ 2.1 GHz, single-threaded, on rev. 526c94c of mathlib4.
The results are summarized in Figure 2. The new Lean implementation is around 30% slower than the C++ implementation, which we attribute mainly to shortcomings in the Lean compiler and data representation compared to C++. Nevertheless, it is within an order of magnitude and practically usable on large libraries, which we consider a strong sign.
7 Related work & Conclusion
Self-verification of ITP systems has been done before. The most relevant references are [2] for Coq, [9] for HOL Light, [5] for Metamath Zero, and to some extent self-verification also overlaps with bootstrapping theorem provers such as Milawa [6] and CakeML [10]. But the most similar work is unquestionably MetaCoq [11], which is a project to develop a verified type-checker for Coq, in Coq. It is significantly more complete than the previous effort [2], but is not yet capable of verifying the Coq standard library because it does not support some of the more unusual or sketchy (mis)features in Coq and there is some effort to get these features removed rather than attempt to validate them. With Lean4Lean we took the approach of quickly getting to feature parity with the real Lean kernel and proving correctness later, so we ended up with a complete verifier but are still lacking in some theorems.
@@ 7 Related work & Conclusion
  • In MetaCoq, the conversion relation is a partial order rather than an equivalence relation, because of universe cumulativity, so 2.13 doesn’t hold as stated. Nevertheless, there are theorems regarding “principal types” which can be used to play the same role.
@@
In this paper we presented a new external verifier for Lean, but it is also the beginning of a larger project to both verify the correctness of a verifier with respect to the Lean metatheory, and also to verify that the Lean metatheory is consistent relative to well-studied axiomatic foundations like ZFC with inaccessible cardinals. Lean 4 is written in Lean, so this is a prime opportunity to have a production-grade ITP which is able to bootstrap at the logical level.

### v2

@@ 6. Results
This is the same interface (and to some extent, the same code) as lean4checker, [^11] https://github.com/leanprover/lean4checker which does the exact same thing but calls addDecl instead, effectively using the C++ kernel as an external verifier for Lean proofs. It may seem odd to use the Lean kernel as an external verifier for itself, but this can catch cases where malicious (or confused) Lean code makes use of the very powerful metaprogramming framework to simply bypass the kernel and add constants to the environment without typechecking them.
@@
lean4checker lean4lean ratio
@@
Batteries 32.49 s 45.74 s 1.40
Mathlib (+ B. + Lean) 44.54 min 58.79 min 1.32
Figure 2. Comparison of the original C++ implementation of the Lean kernel (lean4checker) with lean4lean (Lean) on major Lean packages. Tests were performed on a 12 core 12th Gen Intel i7-1255U @ 2.1 GHz, single-threaded, on rev. 526c94c of mathlib4.
The results are summarized in Figure 2. The new Lean implementation is around 30% slower than the C++ implementation, which we attribute mainly to shortcomings in the Lean compiler and data representation compared to C++. Nevertheless, it is within an order of magnitude and practically usable on large libraries, which we consider a strong sign. We do not expect this to get any faster without Lean compiler improvements, but we also intend to keep the speed from regressing as we add more verification. Because the verification is external to the executable itself, this is possible in theory, but we may need to create new verification tooling to make it doable in practice.
7. Related work & Conclusion
Self-verification of ITP systems has been done in Coq (Barras and Werner, 1997), HOL Light (Harrison, 2006), Isabelle (Nipkow and Roßkopf, 2021), and Metamath Zero (Carneiro, 2020), and to some extent self-verification also overlaps with bootstrapping theorem provers such as Milawa (Davis, 2009) and Candle (Abrahamsson et al., 2022; Kumar et al., 2014). But the most similar work is unquestionably MetaCoq (Sozeau et al., 2020, 2019, 2023), which is a project to develop a verified typechecker for Coq, in Coq. It is significantly more complete than the previous effort (Barras and Werner, 1997), but is not yet capable of verifying the Coq standard library due to lack of eta-conversion and template polymorphism; eta-conversion is work in progress (Lennon-Bertrand, 2022b), and universe polymorphism is planned to be reorganised in Coq’s kernel. With Lean4Lean we took the approach of quickly getting to feature parity with the real Lean kernel and proving correctness later, so we ended up with a feature-complete typechecker but are still lacking in some theorems. Our work also overlaps to some extent with Template-Coq (Anand et al., 2018; Boulier, 2018; Winterhalter, 2020), a certified metaprogramming framework for Coq developed in the MetaCoq repo.
@@ 7. Related work & Conclusion
  • In MetaCoq, the conversion relation is a partial order rather than an equivalence relation, because of universe cumulativity, so Conjecture 2.7 doesn’t hold as stated. Nevertheless, there are theorems regarding “principal types” which can be used to play the same role.
@@
  • Lean’s typechecker is known to not be complete, owing to some of the undecidability results in (Carneiro, 2019), but MetaCoq (Sozeau et al., 2023; Lennon-Bertrand, 2022a) has a proof of completeness.
In this paper we presented a new external typechecker for Lean, but it is also the beginning of a larger project to both verify the correctness of a typechecker with respect to the Lean metatheory, and also to verify that the Lean metatheory is consistent relative to well-studied axiomatic foundations like ZFC with inaccessible cardinals. Lean 4 is written in Lean, so this is a prime opportunity to have a production-grade ITP which is able to bootstrap at the logical level.
Acknowledgements.
I would like to thank Jeremy Avigad for their support and encouragement, Yannick Forster and the anonymous reviewers for reading early drafts of this work, and Matthieu Sozeau and the rest of the MetaCoq team for collaborating and taking interest in this work. Work supported by the Hoskinson Center for Formal Mathematics at Carnegie Mellon.